# The appearance of de Sitter black holes and strong cosmic censorship

### 27 Jan 2024 v1

@@ I Introduction
It was commonly believed that the SCCC holds true and may only be violated in some peculiar modified gravity or when extraordinary matter is present. However, even in the context of GR, which is the most well-established theory of gravity, counterexamples exist. One such counterexample is the RN-de Sitter (dS) black hole, which is distinguished by a single additional parameter compared to the RN black hole. It is one of the most typical solutions with a non-vanishing cosmological constant. Under certain parameters, the RN-dS black hole can exhibit a stable Cauchy horizon Cardoso:2017soq ; Costa:2015pwa ; Costa:2016afl ; Brady:1992cz ; Mellor:1989ac . Consequently, the Penrose diagram of the RN-dS black hole can be extended to be consisted of countless repeated identical universes. Therefore, observers could travel through a black-white hole bridge and then enter another separate universe, which violates the SCCC. More precisely, the Cauchy horizon remains stable when the surface gravity of the Cauchy horizon is less than that of the cosmological horizon Brady:1992 ; Cai:1995ux . The mass inflation in more sophisticate Einstein-Maxwell-scalar field model has been thoroughly studied in the trilogy Costa:2014yha ; Costa:2014zha ; Costa:2014aia , in particular, a detailed range of parameters for mass inflation has been achieved.
@@
Several years ago, the images of the supermassive objects at the center of M87 EventHorizonTelescope:2019dse and Milky Way galaxies EventHorizonTelescope:2022wkp was captured by the Event Horizon Telescope (EHT). This provides us a novel method to detect the compact objects and the black holes. In these black hole pictures, a bright ring surrounds a dark shadow, formed by light emitted directly from the accretion disk. According to calculations for the null geodesics, light rays emitted from the vicinity of the photon sphere, where the effective potential reaches its maximum, can circulate around the black hole numerous times before being received by the observer. However, despite anticipating them, this type of light remained undetected on these images. Utilizing ray-tracing method Gralla:2019xty , we can trace light rays backwards from the observer’s image plane. The trajectory of light rays intersects with the accretion disk, subsequently converging to a bright ring in the image plane. These rings is called a “lensed ring” for two intersections, and “photon ring” for more intersections. However, the Lyapunov exponent being positive indicates that the photon sphere is chaotic. This means that the separation between two initially close trajectories will diverge exponentially over time. Consequently, the photon ring is so close to the lensed ring that it can not be distinguished apart, and is much narrower that its contribution to the overall intensity is negligible. The lensed rings are also covered by the directly emitted light due to the current limitations of astronomical observation precision. As the impact parameter reaches its critical limit, the rays rotate at the photon sphere for infinite times. Since the accretion disk is distributed in a limited range, the rays emitted from the edge of the source will form a sharp edge on the image. The position of the lensed ring or photon ring reveals the information about the background geometry, while the total appearance of a compact object is heavily influenced by the position and profile of the light source.
@@
As the regular black hole with a stable Cauchy horizon, the stable Cauchy horizon in RN-dS black hole also gives rise to the multi-rings structure. Hence it may be a potential method to test SCCC in astronomical observations. However, the multi-ring structure also occurs in the compact object without horizons and the wormholes Olmo:2021piq ; Guerrero:2022qkh ; Guerrero:2022msp ; Olmo:2023lil . The difference between these objects and the black hole violating SCCC is whether they have horizons. In Morales-Herrera:2024zbr , a general relativistic formalism is proposed to get the parameter of the RN black hole (such as the mass and the charge) from some directly observable quantities (such as the total frequency shift, aperture angle of the telescope, and redshift rapidity). Maybe we can distinguish the compact object, wormhole and the RN-dS black hole with the assistance of multiple methods, especially the experiment involving a significant redshift in the presence of an event horizon.
@@ II The ISCO and OSCO in RN-dS black hole
The existence of an ISCO for massive particles in the Schwarzschild black hole at $r=6m$ is a well-known fact. However, in addition to the ISCO, there is an OSCO in the spacetime of the Schwarzschild-dS black hole or more general RN-dS black hole. In this section, we will give a brief introduction to the ISCO and OSCO of the RN-dS black hole.
@@
The stable Cauchy horizon emerges when $\kappa_{-}<\kappa_{\mathrm{c}}$ Brady:1992 ; Cai:1995ux , where $\kappa_{-}$ and $\kappa_{\mathrm{c}}$ are the surface gravities of the Cauchy horizon and cosmological horizon respectively. Once a stable Cauchy horizon exists, the SCCC will be broken down and the predictability of classical theory is threatened. Thus it aroused great interest in the community of general relativity and gravity theory. We will look for the range of the parameters of the black hole which has a stable Cauchy horizon as well as stable circular orbits (i.e., ISCO and OSCO). Firstly, the range of the parameters for the black hole with three horizons has been drawn in Fig.5. A part of the boundary of this range is given by $r_{+}=r_{-}$ and $r_{+}=r_{\mathrm{c}}$. Secondly, the range of the parameters of the RN-dS black hole with a stable Cauchy horizon is shown as the small crescent in Fig.5. The left boundary of this area is $\kappa_{-}=\kappa_{\mathrm{c}}$ and the right boundary is $\kappa_{-}=0$, i.e., $r_{-}=r_{+}$.
@@ III The appearance of the Schwarzschild-dS black hole viewed by observers with different distance
Figure 9: The trajectory of photons. The parameter of black hole is $\Lambda=0.001$. The blue ring is the event horizon and the black ring is the cosmological horizon. The observer is located at $A$. The rays with small $b$ are received by the observer, which are shown as red curves. But some rays with large $b$, whose radius of the perihelion is larger than $r_{\mathrm{obs}}$, can not pass through the observer. This kind of the rays is shown as the orange curve. And the critical case, when the perihelion of the ray is the position of the observer, is shown as the green curve.
The rays with large $b$, like the orange curve, can not be received by the observer because its perihelion is farther than the position of the observer. Therefore, there is a jump in the observed intensity. And the image of the black hole has a significant edge if the observer is close enough to the black hole. In order to avoid this kind of jump, the observer has to be located far way from the black hole. In this case, only light rays with a significantly large value of $b$ cannot be received by the observer. Such light can only be emitted from the remote regions of the accretion disk. Consequently, its contribution to the total intensity is negligible.
@@
Figure 10: Two trajectories of rays from the accretion disk to the observers $A$ at $r_{A}=30$ and $B$ at $r_{B}=48$. The three rings is the Cauchy horizon, event horizon and cosmological horizon of the Schwarzschild-dS black hole with $\Lambda=0.001$ in turn.
@@ IV The appearance of the RN-dS black hole
Figure 12: The Penrose diagram of the RN-dS black hole. The blue curves are the accretion disks in universe $A$ and $B$, respectively. The observer is located near the cosmological horizon. The purple curve is the ray from the accretion disks in universe $B$ and received by the observer. And the red curve is the ray from the accretion disks in universe $A$.
If the Cauchy horizon is stable, the infinite repetition of the spacetime allow the ray emitted from the accretion disk in the previous universe $A$ to fall into the black hole, fly out from the white hole and be received by the observer in universe $B$ finally. This kind of trajectory of the photon is shown as the red curve in Fig.12. Besides, the photons can be emitted from the accretion disk and received by the observer both in universe $B$, as the purple curve in Fig.12. The images caused by this two kinds of ray are quite different. This can be found in the following discussions.
@@ IV.2 The appearance of the RN-dS black hole without ISCO and OSCO
Figure 15: The trajectory of rays around the RN-dS black hole. The black ring is the event horizon. The black line is the the accretion disk between ISCO and OSCO, and the dashed black line is the accretion disk between $r_{+}$ and $r_{\mathrm{c}}$. The rays are for $n=7/4$ (the three rays below) and $n=9/4$ (the three rays above).
The red, purple, and green rays are for the third peak (ring) with the normalized number $n=7/4$ in Fig.14. And the orange, yellow and brown rays are for the fourth one with $n=9/4$. If the accretion disk is between ISCO and OSCO (black line), then the third ring is from $b_{\mathrm{purple}}$ to $b_{\mathrm{green}}$, where $b_{\mathrm{color}}$ is the impact parameter of the ray with corresponding color. And the fourth ring is from $b_{\mathrm{yellow}}$ to $b_{\mathrm{brown}}$. However, if the accretion disk is distribute outside the event horizon (dashed line), then the third ring begins at $b_{\mathrm{red}}$ and the fourth ring begins at $b_{\mathrm{orange}}$. Therefore, the rings caused by the accretion disk outside all the event horizon is wider than the rings caused by the the accretion disk between ISCO and OSCO. Take the yellow ray for an example, it intersects with the accretion disk below twice and the accretion disk above twice. It contributes to the third ring and the fourth ring, too. Thus, all six rays intersect with the accretion disk below for $n=7/4$, and only the orange, yellow and brown rays intersect with accretion disk above for $n=9/4$. This means the fourth ring begins when the third ring is not finished. Therefore, the third and the fourth ring in Fig.14 have a little bit of overlap.
@@ V Conclusions and discussion
SCCC is an important subject in general relativity. It guarantees the predictability of the gravitation theory. However, testing SCCC from astronomical observations has been a challenge. Hence it is significant to study the observational effect of SCCC. The Cauchy horizon is the boundary of the region where the initial data can predict. When the Cauchy horizon is stable, one can pass it and arrive at a area that is not determined by the initial data which violates the SCCC. RN-dS black hole is one of the most famous black hole with a stable Cauchy horizon.

### 02 Mar 2024 v2

@@ I Introduction
It was commonly believed that the SCCC holds true and may only be violated in some peculiar modified gravity or when extraordinary matter is present. However, even in the context of GR, which is the most well-established theory of gravity, counterexamples exist. One such counterexample is the RN-de Sitter (dS) black hole, which is distinguished by a single additional parameter compared to the RN black hole. It is one of the most typical solutions with a non-vanishing cosmological constant. Under certain parameters, the RN-dS black hole can exhibit a stable Cauchy horizon Cardoso:2017soq ; Costa:2015pwa ; Costa:2016afl ; Brady:1992cz ; Mellor:1989ac . Consequently, the Penrose diagram of the RN-dS black hole can be extended to be consisted of countless repeated identical universes. Therefore, observers could travel through a black-white hole bridge and then enter another separate universe, which violates the SCCC. More precisely, the Cauchy horizon remains stable when the surface gravity of the Cauchy horizon is less than that of the cosmological horizon Brady:1992 ; Cai:1995ux . The mass inflation in more sophisticate Einstein-Maxwell-scalar field model has been thoroughly studied in the trilogy Costa:2014yha ; Costa:2014zha ; Costa:2014aia ; Rossetti:2023mrx , in particular, a detailed range of parameters for mass inflation has been achieved.
@@
Several years ago, the images of the supermassive objects at the center of M87 EventHorizonTelescope:2019dse and Milky Way galaxies EventHorizonTelescope:2022wkp was captured by the Event Horizon Telescope (EHT). This provides us a novel method to detect the compact objects and the black holes. In these black hole pictures, a bright ring surrounds a dark shadow, formed by light emitted directly from the accretion disk. According to calculations for the null geodesics, light rays emitted from the vicinity of the photon sphere, where the effective potential reaches its maximum, can circulate around the black hole numerous times before being received by the observer. A more general definition of the photon sphere (surface) can be found in Claudel:2000yi . However, despite anticipating them, this type of light remained undetected on these images. Utilizing ray-tracing method Gralla:2019xty , we can trace light rays backwards from the observer’s image plane. The trajectory of light rays intersects with the accretion disk, subsequently converging to a bright ring in the image plane. These rings is called a “lensed ring” for two intersections, and “photon ring” for more intersections. However, the Lyapunov exponent being positive indicates that the photon sphere is chaotic. This means that the separation between two initially close trajectories will diverge exponentially over time. Consequently, the photon ring is so close to the lensed ring that it can not be distinguished apart, and is much narrower that its contribution to the overall intensity is negligible. The lensed rings are also covered by the directly emitted light due to the current limitations of astronomical observation precision. As the impact parameter reaches its critical limit, the rays rotate at the photon sphere for infinite times. Since the accretion disk is distributed in a limited range, the rays emitted from the edge of the source will form a sharp edge on the image. The position of the lensed ring or photon ring reveals the information about the background geometry, while the total appearance of a compact object is heavily influenced by the position and profile of the light source.
@@
As the regular black hole with a stable Cauchy horizon, the stable Cauchy horizon in RN-dS black hole also gives rise to the multi-rings structure. Hence it may be a potential method to test SCCC in astronomical observations. However, the multi-rings structure also occurs in the compact object without horizons and the wormholes Olmo:2021piq ; Guerrero:2022qkh ; Guerrero:2022msp ; Olmo:2023lil . And the similar multi images structure can be found in the image of stars caused by gravitational lensing of a massive object Tsukamoto:2021caq ; Tsukamoto:2022vkt . The difference between these objects and the black hole violating SCCC is whether they have horizons. In Morales-Herrera:2024zbr , a general relativistic formalism is proposed to get the parameter of the RN black hole (such as the mass and the charge) from some directly observable quantities (such as the total frequency shift, aperture angle of the telescope, and redshift rapidity). And a method to get the parameter of the Kerr-dS black hole is proposed in Momennia:2023lau . Maybe we can distinguish the compact object, wormhole and the RN-dS black hole with the assistance of multiple methods, especially the experiment involving a significant redshift in the presence of an event horizon.
@@ II The ISCO and OSCO in RN-dS black hole
The existence of an ISCO for massive particles in the Schwarzschild black hole at $r=6m$ is a well-known fact. However, in addition to the ISCO, there is an OSCO in the spacetime of the Schwarzschild-dS black hole or more general RN-dS black hole. Physically acceptable substances that can emit light are generally massive. As a result, we adapt the model of the accretion disk, which is regarded as the source of light emission, to consist of the stable orbits of these massive substances. Consequently, the accretion disk in an RN-dS black hole resembles a band distributed between the ISCO and OSCO. The accretion disk of a Kerr-dS spacetime is studied in Stuchlik:2020rls . In this section, we will give a brief introduction to the ISCO and OSCO of the RN-dS black hole.
@@
The stable Cauchy horizon emerges when $\kappa_{-}<\kappa_{\mathrm{c}}$ Brady:1992 ; Cai:1995ux , where $\kappa_{-}$ and $\kappa_{\mathrm{c}}$ are the surface gravities of the Cauchy horizon and cosmological horizon respectively. Once a stable Cauchy horizon exists, the SCCC will be broken down and the predictability of classical theory is threatened. Thus it aroused great interest in the community of general relativity and gravity theory. We will look for the range of the parameters of the black hole which has a stable Cauchy horizon as well as stable circular orbits (i.e., ISCO and OSCO). Firstly, the range of the parameters for the black hole with three horizons has been drawn in Fig.5. A part of the boundary of this range is given by $r_{+}=r_{-}$ and $r_{+}=r_{\mathrm{c}}$. The parameter range for Kerr-dS black holes can be found in Momennia:2023lau . Although their graph closely resembles ours, it is important to note that their abscissa represents the angular momentum parameter $\alpha$, while our abscissa corresponds to the charge parameter $Q$. This observation highlights the similarity between angular momentum and charge, illustrating that the RN-dS black hole can serve as a valuable reference for further study on Kerr-dS black holes. Secondly, the range of the parameters of the RN-dS black hole with a stable Cauchy horizon is shown as the small crescent in Fig.5. The left boundary of this area is $\kappa_{-}=\kappa_{\mathrm{c}}$ and the right boundary is $\kappa_{-}=0$, i.e., $r_{-}=r_{+}$.
@@ III The appearance of the Schwarzschild-dS black hole viewed by observers with different distance
Figure 9: The trajectory of photons. The parameter of black hole is $\Lambda=0.001$. The blue ring is the event horizon and the black ring is the cosmological horizon. The observer is located at $A$. The rays with small $b$ are received by the observer, which are shown as red curves. But some rays with large $b$, whose radius of the perihelion is larger than $r_{\mathrm{obs}}$, can not pass through the observer. This kind of the rays is shown as the blue curve. And the critical case, when the perihelion of the ray is the position of the observer, is shown as the green
curve.
Figure 9: The trajectory of photons. The parameter of black hole is $\Lambda=0.001$. The blue ring is the event horizon and the black ring is the cosmological horizon. The observer is located at $A$. The rays with small $b$ are received by the observer, which are shown as red curves. But some rays with large $b$, whose radius of the perihelion is larger than $r_{\mathrm{obs}}$, can not pass through the observer. This kind of the rays is shown as the blue curve. And the critical case, when the perihelion of the ray is the position of the observer, is shown as the green
The rays with large $b$, like the blue curve, can not be received by the observer because its perihelion is farther than the position of the observer. Therefore, there is a jump in the observed intensity. And the image of the black hole has a significant edge if the observer is close enough to the black hole. In order to avoid this kind of jump, the observer has to be located far way from the black hole. In this case, only light rays with a significantly large value of $b$ cannot be received by the observer. Such light can only be emitted from the remote regions of the accretion disk. Consequently, its contribution to the total intensity is negligible.
@@
Figure 10: Two trajectories of rays from the accretion disk to the observers $A$ at $r_{A}=30$ and $B$ at $r_{B}=48$. The three rings are the Cauchy horizon, event horizon and cosmological horizon of the Schwarzschild-dS black hole with $\Lambda=0.001$ in turn.
@@ IV The appearance of the RN-dS black hole
Figure 12: The Penrose diagram of the RN-dS black hole. The blue curves are the accretion disks in universe $A$ and $B$, respectively. The observer is located near the cosmological horizon. The green curve is the ray from the accretion disks in universe $B$ and received by the observer. And the red curve is the ray from the accretion disks in universe $A$.
If the Cauchy horizon is stable, the infinite repetition of the spacetime allow the ray emitted from the accretion disk in the previous universe $A$ to fall into the black hole, fly out from the white hole and be received by the observer in universe $B$ finally. This kind of trajectory of the photon is shown as the red curve in Fig.12. Besides, the photons can be emitted from the accretion disk and received by the observer both in universe $B$, as the green curve in Fig.12. The images caused by this two kinds of ray are quite different. This can be found in the following discussions.
@@ IV.2 The appearance of the RN-dS black hole without ISCO and OSCO
Figure 15: The trajectory of rays around the RN-dS black hole. The black ring is the event horizon. The black line is the the accretion disk between ISCO and OSCO, and the dashed black line is the accretion disk between $r_{+}$ and $r_{\mathrm{c}}$. The rays are for $n=7/4$ (the three solid rays labeled 1$\sim$3 in (a)) and $n=9/4$ (the three dashed rays labeled 4$\sim$6 in (b) )
Figure 15: The trajectory of rays around the RN-dS black hole. The black ring is the event horizon. The black line is the the accretion disk between ISCO and OSCO, and the dashed black line is the accretion disk between $r_{+}$ and $r_{\mathrm{c}}$. The rays are for $n=7/4$ (the three solid rays labeled 1$\sim$3 in (a)) and $n=9/4$ (the three dashed rays labeled 4$\sim$6 in (b) )
The red, green, and blue solid rays are for the third peak (ring) with the normalized number $n=7/4$ in Fig.14. And the red, green, and blue dashed rays are for the fourth one with $n=9/4$. If the accretion disk is between ISCO and OSCO (black line), then the third ring is from $b_{2}$ to $b_{3}$, where $b_{i}$ is the impact parameter of the ray with corresponding number labeled in Fig. 15. And the fourth ring is from $b_{5}$ to $b_{6}$. However, if the accretion disk is distribute outside the event horizon (black dashed line), then the third ring begins at $b_{1}$ and the fourth ring begins at $b_{4}$. Therefore, the rings caused by the accretion disk outside all the event horizon is wider than the rings caused by the the accretion disk between ISCO and OSCO. Take the green dashed ray (labeled 5) for an example, it intersects with the accretion disk below twice and the accretion disk above twice. It contributes to the third ring and the fourth ring, too. Thus, all six rays intersect with the accretion disk below for $n=7/4$, and only the red, green, and blue dashed rays intersect with accretion disk above for $n=9/4$. This means the fourth ring begins when the third ring is not finished. Therefore, the third and the fourth ring in Fig.14 have a little bit of overlap.
@@ V Conclusions and discussion
SCCC is an important subject in general relativity. It guarantees the predictability of the gravitation theory. However, although some potential observations violating WCCC is proposed in the RN singularity Wagner:2023axm , testing SCCC from astronomical observations has been a challenge. Hence it is significant to study the observational effect of SCCC. The Cauchy horizon is the boundary of the region where the initial data can predict. When the Cauchy horizon is stable, one can pass it and arrive at a area that is not determined by the initial data which violates the SCCC. RN-dS black hole is one of the most famous black hole with a stable Cauchy horizon.
@@
Within classical theory, some black holes have a stable Cauchy horizons. However, certain studies suggest that in such scenarios, quantum effects will play an important role and eventually render these Cauchy horizons unstable Cai:1995nt ; Cai:1998yp ; Hollands:2019whz . Nevertheless, whether quantum effects will overturn established classical results remains an unresolved issue. Probing the multi-ring structure could help examine the SCCC, and further assessing the significance of quantum effects.